# Unveiling Challenges for LLMs in Enterprise Data Engineering

### v1

@@ Abstract.
Large Language Models (LLMs) have demonstrated significant potential for automating data engineering tasks on tabular data, giving enterprises a valuable opportunity to reduce the high costs associated with manual data handling. However, the enterprise domain introduces unique challenges that existing LLM-based approaches for data engineering often overlook, such as large table sizes, more complex tasks, and the need for internal knowledge. To bridge these gaps, we identify key enterprise-specific challenges related to data, tasks, and background knowledge and conduct a comprehensive study of their impact on recent LLMs for data engineering. Our analysis reveals that LLMs face substantial limitations in real-world enterprise scenarios, resulting in significant accuracy drops. Our findings contribute to a systematic understanding of LLMs for enterprise data engineering to support their adoption in industry.
@@ 1. Introduction
Large enterprises generate vast amounts of tabular data that drives applications like machine learning and analytical query processing. Data engineering is crucial for understanding this raw data and transforming it into a suitable form for its downstream usage. It encompasses a range of tasks, from data exploration and transformation to data integration and cleaning. Since such tasks often impose significant manual overhead to apply existing tools to the specific data at hand, the automation of individual data engineering tasks like entity matching (Konda et al., 2016; Mudgal et al., 2018) and column type annotation (Hulsebos et al., 2019; Zhang et al., 2020a) with the help of machine learning has long drawn attention from researchers. Nevertheless, adapting such machine learning approaches to new datasets and tasks often requires computer science expertise, rendering them inaccessible to many practitioners.
LLMs to the rescue?
@@
Recent work has shown that Large Language Models (LLMs) such as GPT-4 (OpenAI et al., 2024) can be directly applied to data engineering tasks on tabular data, indicating that they achieve state-of-the-art results on various table-based tasks without requiring task-specific architectures and training (Narayan et al., 2022; Jaimovitch-López et al., 2022; Kayali et al., 2024; Barke et al., 2024). Their out-of-the-box nature is a significant advantage over other machine learning approaches that require supervised training for each dataset and task. One example relevant for data understanding is the task of column type annotation, where the goal is to annotate the columns of a relational table with semantic types from a given ontology. Whereas machine learning-based approaches like Sherlock (Hulsebos et al., 2019) and Sato (Zhang et al., 2020a) require re-training for each new set of semantic types, LLMs can easily support different sets of semantic types by including them in the prompt (Korini and Bizer, 2023). Therefore, they could provide a promising avenue to automate data engineering tasks.
Is this also true for enterprise data?
@@
Recently, some first papers have made the point that while LLMs achieve good results on existing benchmarks, they fail miserably when applied to real-world enterprise data (Kayali et al., 2025; Bodensohn et al., 2024b). A central observation here is that the data in public benchmarks is often crawled from web resources like Wikipedia (Bhagavatula et al., 2015) and GitHub (Hulsebos et al., 2023). In contrast, the data from companies running their business processes with software systems like those from SAP differs fundamentally from these datasets in many aspects, including table sizes, sparsity, and data types (Kayali et al., 2025; Savelieva et al., 2022; Vogelsgesang et al., 2018). Since LLMs are typically trained on public data scraped from the web (Brown et al., 2020; Meta, 2024), they have not seen significant amounts of such enterprise data during their training.
A significant performance decline on enterprise data.
@@
In this paper, we thus set out the goal of systematically studying the performance of LLMs for enterprise data engineering. As a “sneak peak” into our results, we demonstrate the results of a first experiment on enterprise data where we compare the performance of LLMs on the task of column type annotation on a public benchmark to a dataset of real-world customer data from SAP. As shown in Figure 1 (left bar group), LLMs of different types and generations achieve high F1 scores of up to 0.94 when solving this task on a public corpus like the SportsTables dataset (Langenecker et al., 2023). In contrast, as shown in Figure 1 (second bar group), using LLMs on the representative customer data from SAP for column type annotation with a comparable number of semantic types causes a substantial performance decrease across all models, ranging from single-digit F1 scores of 0.07 up to 0.38.
Beyond data challenges.
@@
Beyond these data challenges, data engineering in enterprises faces additional difficulties:
@@
(1) Task complexity: Enterprise tasks are often more complex than their academic formulations (Savelieva et al., 2022). For example, literature about the task of entity matching often assumes that each entity is one row in a single table (Konda et al., 2016; Peeters and Bizer, 2023). By contrast, entities in enterprises are often business objects that span across multiple tables, making matching notoriously more difficult. Moreover, the tasks themselves are often more complex. For example, increasing the task complexity of column type annotation by scaling to the true number of semantic types (5,089 instead of 200) in the SAP system leads to a further 50% drop in F1 scores, as shown in Figure 1 (+ Task Challenges).
@@
(2) Internal knowledge: Data engineering in enterprises also often requires internal knowledge that is absent from public sources, limiting LLMs’ capabilities to understand the data without additional knowledge. This is especially true for schema customizations, which involve customer-defined semantic types. On SAP’s column type annotation dataset, when adding those columns to the task, F1 scores for such customer-defined columns are near zero, as shown in Figure 1 (+ Knowledge Challenges).
A broad analysis of enterprise data engineering.
In this paper, we conduct a broad experimental analysis on representative customer data and case studies reflecting real enterprise scenarios. Our goal is to unmask the jagged out-of-the-box capabilities of LLMs to provide insights into where they can be reliably used and where additional tweaks and improvements are necessary. To accomplish this, we apply five LLMs from three model providers to various established data engineering tasks. As the main goal of this paper, we want to highlight how particular enterprise-specific challenges impact LLM performance and provide insights that can guide future efforts to make LLMs viable for enterprise data engineering.
We believe this to be the first attempt to examine LLMs for data engineering on real enterprise data at this breadth. We see it as an important first step (of many) to make LLMs viable for enterprise data engineering.
@@
Our key contributions in this paper are:111This paper extends our previously published work (Bodensohn et al., 2024b, a) on this topic. (1) We systematically analyze the challenges involved in enterprise data engineering and structure them into aspects regarding enterprise data, enterprise tasks, and enterprise knowledge. (2) We experiment on representative enterprise data to show how it differs from existing public benchmarks and understand its impact on LLM performance. (3) We conduct multiple case studies that reflect real-world enterprise scenarios, allowing us to extensively evaluate how different challenges affect LLM performance in isolation. (4) We discuss potential directions for addressing these challenges, as well as the costs of using LLMs at enterprise scale. (5) Our experiments are performed with five recent LLMs from three model providers (OpenAI, Anthropic, and Meta). To enable follow-up research, we make the code—including our full evaluation setup, all prompts, and where legally possible also our data—of this paper available to the broader research community.
Outline.
The paper is structured as follows: Section 2 discusses the setting of our study on LLMs for data engineering, and Section 3 introduces our methodology for analyzing the challenges of using LLMs in enterprises. Sections 4 through 6 then present the analysis, experimental results, and discussions for the identified categories of challenges: data, tasks, and knowledge. Section 7 discusses the costs involved in enterprise-scale data engineering. Finally, Section 8 summarizes our conclusions and avenues for future research.
@@
Pricing by OpenAI, Anthropic, and AWS Bedrock3 in Feb 2025.
@@ 2. Setting of our Study
Using LLMs to solve table-based tasks is a promising research direction that has been actively studied in recent years (Narayan et al., 2022; Zhang et al., 2024a; Peeters and Bizer, 2023; Korini and Bizer, 2023; Zhang et al., 2024b). This section introduces the models we evaluate and briefly summarizes existing research on LLMs for data engineering.
@@
LLMs are text foundation models trained on large corpora to complete natural language inputs (Brown et al., 2020) and follow user instructions (Wei et al., 2022; Liu et al., 2023). Modern LLMs like GPT-4 (OpenAI et al., 2024) and Llama 3 (Grattafiori et al., 2024) build on a range of further ideas, from mixture-of-experts architectures (Shazeer et al., 2017) to more efficient implementations (Dao et al., 2022), leading to continuous improvements in abilities like applying background knowledge and handling long inputs. With the release of reasoning models such as OpenAI’s o1 (OpenAI, 2024b) and Deepseek’s R1 (DeepSeek-AI et al., 2025), LLMs perform even better at tasks that require several intermediate steps. While many LLMs would be relevant candidates to include in our evaluation, our selection is necessarily limited. To ensure generalizable results, we evaluate five recent LLMs from three model providers covering multiple sizes and costs and including open and closed models. As shown in Table 1, we use GPT-4o (OpenAI, 2024a) as the state-of-the-art and GPT-4o-Mini as the cheapest model from OpenAI as well as o1 (OpenAI, 2024c) as a reasoning model. We further include Claude 3.5 Sonnet (Anthropic, 2024) from Anthropic and Llama 3.1 Instruct (Meta, 2024) from Meta in our evaluations.
Positioning this study.
Although several papers have already observed the differences between enterprise and web data (Vogelsgesang et al., 2018; Savelieva et al., 2022; Kang and Naughton, 2003; Zhu et al., 2019), existing research on LLMs for data engineering primarily uses evaluation datasets based on tables from public web sources, calling the applicability of LLMs on real-world enterprise data into question. More recently, some enterprise-specific benchmarks have been released for tabular prediction tasks (Klein et al., 2024), Text-to-SQL (Chen et al., 2024), and column type annotation (Kayali et al., 2025). Related to our work, in a recent study, Kayali et al. (2025) quantify performance gaps between private and public data for the task of semantic column type annotation and find that benchmarks based on public data overestimate the performance of LLMs. However, these studies focus only on individual tasks and consider the data as the only difference in enterprise scenarios. In contrast, we perform case studies on various diverse tasks to systematically analyze the challenges along several dimensions (data, tasks, and knowledge) of working in enterprise scenarios.
@@ 3. Design of our Study
Non-numeric (abc) and numeric (123) columns determined by pandas.
@@
We aim to take a holistic view of data engineering in enterprises by systematically analyzing the out-of-the-box performance of LLMs, covering enterprise-specific challenges along the dimensions of data, tasks, and knowledge. Figure 2 provides an overview of the tasks we have chosen for analyzing the performance along each dimension. Below, we explain the rationale for selecting these tasks as examples to examine challenges that arise broadly across different enterprise tasks and scenarios. Sections 4 to 6 then describe our experimental evaluation for each of these challenges.
@@
Enterprise data differs from public evaluation datasets in various aspects. First, tables in enterprise databases are substantially larger in their number of columns and rows, and the schema and data are often more complex with table names, column names, and values that lack intuitive meaning. Moreover, enterprise data exhibits a much higher sparsity compared to public data, leaving many cells empty. Together, these factors make data engineering on enterprise data much more challenging. To explore these challenges, in our study, we focus on the task of column type annotation (see Figure 2 left) (Hulsebos et al., 2019; Zhang et al., 2020a). We have chosen this task because it is well-studied in the literature and thus enables the comparison of accuracies with public datasets. Moreover, the task can be formulated in various alternative ways, stressing different data factors (e.g., with and without schema information as input), allowing us to study the effects of different challenges like table size, sparsity, and descriptiveness on the data and schema level.
@@
Beyond the complexities of enterprise data, enterprise tasks themselves are also more complex. First, while academia typically studies data engineering tasks in isolation, enterprise tasks in practice are often compounds of multiple simpler data engineering steps. As such, we argue that the evaluation procedures themselves must change and we need to study accuracy end-to-end and analyze effects such as how errors propagate. In this paper, we examine the compound nature of enterprise tasks using the example of integrating two customer databases, as illustrated in Figure 2 (middle). This process involves the tasks of schema matching, entity matching, and data integration. A second challenge for enterprise tasks (not shown in Figure 2) is that even the individual steps are often more complex. For example, while entity matching in academia assumes 1:1 matches across rows of two tables (Peeters et al., 2024; Mudgal et al., 2018), matching in enterprise scenarios covers more complex scenarios like mapping several bank transfers to one invoice, and the required data is often scattered across multiple tables.
@@
As a last challenge, data engineering in enterprises often requires enterprise-specific knowledge. This is particularly challenging because LLMs will likely not have seen the required information during training since it is covered only in internal documentation or even just certain implementation details in the code of enterprise systems. To analyze such challenges, we have selected a task related to data exploration, which requires translating natural language queries about business processes into the enterprise-specific query language SIGNAL, which differs slightly from SQL (Figure 2 right) (Kampik et al., 2023). The task is interesting because LLMs should, judging by its similarity to Text-to-SQL, in principle be able to solve it. However, they lack particular information about how exactly SIGNAL differs from SQL. At the core, the question we want to answer here is to what extent the lack of enterprise knowledge affects the accuracy of LLMs, and whether it can be increased by providing enterprise-specific knowledge as context to LLMs.
@@
Since enterprise tables are large and can contain millions of rows, using LLMs on such data to solve complex multi-step tasks can cause high costs, rendering some of the larger, more complex LLMs economically unviable. We discuss this aspect in Section 7.

### v2

@@ Abstract.
Large Language Models (LLMs) promise to automate data engineering on tabular data, offering enterprises a valuable opportunity to cut the high costs of manual data handling. But the enterprise domain comes with unique challenges that existing LLM-based approaches for data engineering often overlook, such as large table sizes, more complex tasks, and the need for internal knowledge. To bridge these gaps, we identify key enterprise-specific challenges related to data, tasks, and background knowledge and extensively evaluate how they affect data engineering with LLMs. Our analysis reveals that LLMs face substantial limitations in real-world enterprise scenarios, with accuracy declining sharply. Our findings contribute to a systematic understanding of LLMs for enterprise data engineering to support their adoption in industry.
PVLDB Reference Format:
PVLDB, 19(2): 196 - 209, 2025.
doi:10.14778/3773749.3773758
††This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 19, No. 2 ISSN 2150-8097.
doi:10.14778/3773749.3773758
PVLDB Artifact Availability:
@@ 1. Introduction
Large enterprises generate vast amounts of tabular data that drives applications like machine learning and analytical query processing. Data engineering is crucial for understanding this raw data and preparing it for its downstream usage. It encompasses a range of tasks, from data exploration and integration to data transformation and cleaning. Since these tasks often impose significant manual overhead to apply existing tools to the specific data at hand, the automation of individual data engineering tasks like entity matching (Konda et al., 2016; Mudgal et al., 2018) and column type annotation (Hulsebos et al., 2019; Zhang et al., 2020) with the help of machine learning has long drawn attention from researchers. Nevertheless, adapting such machine learning approaches to new datasets and tasks often requires computer science expertise, rendering them inaccessible to many practitioners.
@@
Recent work has shown that Large Language Models (LLMs) can be directly applied to data engineering tasks on tabular data, indicating that they achieve state-of-the-art results on various table-based tasks without requiring task-specific architectures and training (Narayan et al., 2022; Jaimovitch-López et al., 2022; Kayali et al., 2024; Barke et al., 2024). Their out-of-the-box nature provides a significant advantage over other machine learning approaches that require supervised training for each dataset and task. One example is the task of column type annotation, where the goal is to annotate the columns of a relational table with semantic types from a given ontology. Whereas machine learning-based approaches like Sherlock (Hulsebos et al., 2019) and Sato (Zhang et al., 2020) require re-training for each new set of semantic types, LLMs can easily support different sets of semantic types by including them in the prompt (Korini and Bizer, 2023). Therefore, they provide a promising avenue to automate data engineering tasks.
Enterprise challenges.
@@
Some first papers have now made the point that while LLMs achieve strong results on existing benchmarks, they fall short when applied to real-world enterprise data (Kayali et al., 2025; Bodensohn et al., 2024b). A central observation here is that the data in public benchmarks often comes from web resources like Wikipedia (Bhagavatula et al., 2015) and GitHub (Hulsebos et al., 2023). In contrast, the data from companies running their business processes with software systems like those from SAP differs fundamentally from these datasets in many aspects, including table sizes, sparsity, and data types (Kayali et al., 2025; Savelieva et al., 2022; Vogelsgesang et al., 2018). Since LLMs are typically trained on public data scraped from the web (Brown et al., 2020; Meta, 2024), they have not seen much of this enterprise data during their training.
@@
In this paper, we set out to systematically study the performance of LLMs for enterprise data engineering on representative customer data and case studies reflecting real enterprise scenarios. As a sneak peak into our results, Figure 1 presents a first experiment comparing the performance of LLMs for column type annotation on a public benchmark to a dataset of real-world customer data from SAP. As shown by the left bar group, LLMs of different types and sizes achieve high F1 scores when evaluated on a public benchmark like the SportsTables dataset (Langenecker et al., 2023). The same models, however, show substantial performance decreases when applied to the representative customer data from SAP for column type annotation with a comparable number of semantic types, as shown in Figure 1 (SAPCTA ).
@@
Aside from these data challenges, data engineering in enterprises comes with additional difficulties:
@@
(1) Task complexity: Enterprise tasks are often more complex than their academic formulations (Savelieva et al., 2022). One example is the task of entity matching, where existing literature often assumes that each entity is one row in a single table (Konda et al., 2016; Peeters and Bizer, 2023). In enterprises, however, entities are often business objects that span across multiple tables, making matching difficult. Moreover, the tasks themselves are often more complex. For example, increasing the complexity of column type annotation by scaling to the true number of semantic types (5,089 instead of 200) in the SAP system leads to a further decrease in F1 scores, as shown in Figure 1 (+ Task Challenges).
@@
(2) Internal knowledge: Data engineering in enterprises also often requires internal knowledge that is absent from public sources, limiting LLMs’ abilities to understand the data without additional context. This is especially true for schema customizations that involve customer-defined semantic types. On SAP’s column type annotation dataset, when adding those columns to the task, F1 scores for such customer-defined columns are near zero, as shown in Figure 1 (+ Knowledge Challenges).
We want to highlight how such enterprise-specific challenges affect LLMs. Our findings go beyond observations already known to the community and help unmask the jagged out-of-the-box capabilities of LLMs. For example, the impact of the complex enterprise schemas and symbolic values in the enterprise data is particularly striking. Moreover, we were surprised by how strongly the models resisted overwriting their parametric knowledge despite explicit guidance from documentation and examples. We believe this to be the first attempt to examine LLMs for data engineering on real enterprise data at this breadth. We see it as an important first step (of many) to make LLMs viable for enterprise data engineering.
@@
Our main contributions are:111This paper extends our previously published work (Bodensohn et al., 2024b, a) on this topic. (1) We systematically analyze the challenges involved in enterprise data engineering and structure them along the dimensions of enterprise data, enterprise tasks, and enterprise knowledge. (2) We experiment on representative enterprise data to show how it differs from existing public benchmarks and understand how it affects LLMs. (3) We conduct multiple case studies that reflect real-world enterprise scenarios, allowing us to evaluate how different challenges affect LLMs in isolation. (4) We discuss directions for addressing these challenges, as well as the costs of using LLMs at enterprise scale. (5) Our experiments are performed with five recent LLMs from OpenAI, Anthropic, and Meta. To enable follow-up research, we make the code—including the full evaluation setup, all prompts, and where legally possible also the data—of this paper available to the broader research community.
@@
Pricing by OpenAI, Anthropic, and AWS Bedrock3 in February 2025.
@@ 2. Setting of our Study
Using LLMs to solve table-based tasks is a promising research direction that has been actively studied in recent years (Narayan et al., 2022; Zhang et al., 2024a; Peeters and Bizer, 2023; Korini and Bizer, 2023; Zhang et al., 2024b). This section introduces the models we evaluate and briefly summarizes existing research on data engineering with LLMs.
@@
LLMs are text foundation models trained on large corpora to complete natural language inputs (Brown et al., 2020) and follow user instructions (Wei et al., 2022; Liu et al., 2023). Recent LLMs like GPT-4 (OpenAI et al., 2024) and Llama 3 (Grattafiori et al., 2024) build on a range of further ideas, from mixture-of-experts architectures (Shazeer et al., 2017) to more efficient implementations (Dao et al., 2022), leading to improvements in abilities like applying background knowledge and handling long inputs. With the release of reasoning LLMs like OpenAI’s o1 (OpenAI, 2024b) and Deepseek’s R1 (DeepSeek-AI et al., 2025), they perform even better on tasks requiring several intermediate steps.
While many LLMs would be relevant candidates to include in our evaluation, our selection is necessarily limited. To ensure generalizable results, we evaluate five recent LLMs from three model providers covering multiple types and sizes and including open and closed models. As shown in Table 1, we use GPT-4o (OpenAI, 2024a) as the state-of-the-art and GPT-4o-Mini as the cheapest model from OpenAI in February 2025, and o1 (OpenAI, 2024c) as a reasoning model. We further include Claude 3.5 Sonnet (Anthropic, 2024) from Anthropic and Llama 3.1 Instruct (Meta, 2024) from Meta. In general, LLMs bear the challenge of output variability. In all our experiments, we reduce randomness to a minimum by setting the sampling temperature to 0. We also repeat several of our experiments multiple times, witnessing only small standard errors and no change in the overall characteristics of the results.
Positioning the study.
Although several papers have already observed the differences between enterprise and web data (Vogelsgesang et al., 2018; Savelieva et al., 2022; Kang and Naughton, 2003; Zhu et al., 2019), existing research on data engineering with LLMs primarily uses evaluation datasets based on tables from public web resources, calling the applicability of LLMs on real-world enterprise data into question. More recently, some enterprise-specific benchmarks have been released for tabular prediction tasks (Klein et al., 2024) and Text-to-SQL (Papicchio et al., 2025; Chen et al., 2024). Related to our work, Kayali et al. (2025) quantify performance gaps between private and public data for the task of column type annotation and find that benchmarks based on public data overestimate the performance of LLMs. However, these studies focus only on individual tasks and consider the data as the only challenge in enterprise scenarios. In contrast, we perform case studies on various tasks to systematically analyze the challenges along several dimensions of working in enterprise scenarios.
@@ 3. Design of our Study
Non-numerical (abc) and numerical (123) columns determined by pandas.
@@
We aim to take a holistic view of data engineering in enterprises by systematically analyzing the out-of-the-box performance of LLMs, covering enterprise-specific challenges along the dimensions of data, tasks, and knowledge. Figure 2 provides an overview of the tasks we have chosen to analyze the performance along each dimension. Below, we explain the rationale for selecting these tasks as examples to examine challenges that arise broadly across different enterprise tasks and scenarios. Sections 4 to 6 then describe our experimental evaluation for each of these challenges.
@@
Enterprise data differs from public evaluation datasets in various aspects. First, the tables in enterprise databases are substantially larger in their number of columns and rows, and the schema and data are often more complex with column names and values that lack intuitive meaning. Moreover, enterprise data exhibits a much higher sparsity than public data, leaving many cells empty. Together, these factors make data engineering on enterprise data much more challenging. To explore these challenges, we focus on the task of column type annotation (Figure 2 left) (Hulsebos et al., 2019; Zhang et al., 2020). We have chosen this task because it is well-studied in the literature and thus enables the comparison of accuracies with public datasets. Moreover, the task can be formulated in various ways, stressing different data factors (e.g., with and without schema information) and allowing us to study the effects of different challenges like table size, sparsity, and descriptiveness of the data and schema.
@@
Beyond the complexities of enterprise data, the enterprise tasks themselves are also more complex. First, while academia typically studies data engineering tasks in isolation, enterprise tasks in practice are often compounds of multiple simpler data engineering steps. As such, we argue that the evaluation procedures themselves must change and we need to study accuracy end-to-end to analyze effects such as how errors propagate. In this paper, we examine the compound nature of enterprise tasks using the example of integrating two customer databases, as illustrated in Figure 2 (middle). This process involves the steps of schema matching, entity matching, and data integration. A second challenge for enterprise tasks (not shown in Figure 2) is that even the individual steps are often more complex. For example, while entity matching in academia assumes 1:1 matches across rows of two tables (Peeters et al., 2024; Mudgal et al., 2018), matching in enterprise scenarios covers more complex cases like mapping several bank transfers to one invoice, and the required data is often scattered across multiple tables.
@@
As a last challenge, data engineering in enterprises often requires enterprise-specific knowledge. This is particularly challenging because LLMs will likely not have seen the required information during training, as it is covered only in internal documentation or even just certain implementation details in the code of enterprise systems. To analyze these challenges, we have selected a task related to data exploration that requires translating natural language queries about business processes into the enterprise-specific query language SIGNAL, which differs slightly from SQL (Figure 2 right) (Kampik et al., 2023). The task is interesting because LLMs should, judging by its similarity to Text-to-SQL, in principle be able to solve it. However, they lack particular information about how exactly SIGNAL differs from SQL. At the core, the question we want to answer here is to what extent the lack of enterprise knowledge affects the accuracy of LLMs, and whether it can be increased by providing enterprise-specific knowledge as context.
@@
Since enterprise tables can contain millions of rows, using LLMs on such data can cause high costs, rendering some of the larger, more complex LLMs economically unviable. We discuss this aspect in Section 7.